# Deconfined quantum criticality of frustrated hard-core dipolar bosons

### v2

@@ Abstract
where $\chi_{L}^{m}$ represents the fidelity susceptibility per site at the peak position $\lambda=\lambda_{m}$, which converges towards the critical point $\lambda_{c}$ as $L$ tends to infinity. As the AFM phase breaks the $\mathbb{Z}_{2}$ spin-flip symmetry and the VBS phase breaks translation symmetry, the direct continuous second-order phase transition occurring between the AFM phase and the VBS state indicates that the phase transition is a DQCP beyond the LGW theory.
Interestingly, the DQCP arises from the merging of two BKT transitions. These transitions can be understood through the concept of domain walls, which are topological defects in one-dimensional systems. Beginning from the VBS phase (i.e at $t_{2}/t_{1}=0.7$ and $V=0$), increasing the interaction strength $V/t_{1}$ induces the formation of AFM domain walls. Domain walls in VBS bind the AFM phase (and vice versa), leading to the destruction of the VBS phase and the emergence of the AFM phase as $V/t_{1}$ increases further. Consequently, a direct transition between the VBS phase to the AFM phase occurs under conditions of large $t_{2}/t_{1}$ and $V/t_{1}$, indicating the presence of a DQCP. This DQCP is linked to the two BKT transitions through a multicritical point characterized by an emergent high symmetry.
@@
Conclusion-. In summary, we employed the DMRG method to investigate novel physical phenomena in the frustrated dipolar Bose-Hubbard model and to provide a comprehensive phase diagram of the model through the exploration of fidelity susceptibilities and order parameters. We identified two gapped phases, the AFM phase and the VBS phase, along with one gapless XY phase, each exhibiting two BKT transitions and a DQCP. Remarkably, we found that the DQCP arises from the fusion of two BKT transitions in this model. Moreover, we propose that this model could be realized using dipolar bosons.
Acknowledgments-. G.S. thanks Luis Santos for the valuable comments. G.S. is appreciative of support from the NSFC under the Grants No. 11704186, "the Fundamental Research Funds for the Central Universities, NO. NS2023055". W.-L.Y is supported by the NSFC under Grant No. 12174194, Opening Fund of the Key Laboratory of Aerospace Information Materials and Physics (Nanjing University of Aeronautics and Astronautics), MIIT, Top-notch Academic Programs Project of Jiangsu Higher Education Institutions (TAPP), and stable supports for basic institute research under Grant No. 190101. Y.-N.W and W.-Y.Z are supported by the Postgraduate Research & Practice Innovation Program of Jiangsu Province under Nos. KYCX24_0526 and KYCX23_0347, respectively. S.-P. K is appreciative of support from the NSFC under the Grant Nos. 11974053 and 12174030, and the National Key R&D Program of China under Grant No. 2023YFA1406704. This work is partially supported by the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics.
Supplemental Material
We present additional information on (1) Non-interacting energy spectrum, (2) Level crossing for the BKT transition and (3) Correlation length critical exponent.
A: Non-interacting energy spectrum
@@
$\begin{split}b_{r}=\frac{1}{\sqrt{N}}\sum e^{ikj}b_{k},\\ b_{r}^{\dagger}=\frac{1}{\sqrt{N}}\sum e^{-ik^{\prime}j}b_{k^{\prime}}^{\dagger},\end{split}$ (S2)
the model described in Eq.(S1) can be transformed into momentum space,
$\begin{split}H&=t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{2ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-2ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}\\ &=\sum_{k}\left[t_{1}(e^{ik}+e^{-ik})+t_{2}(e^{2ik}+e^{-2ik})\right]b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\left(2t_{1}\cos{k}+2t_{2}\cos{2k}\right)b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\epsilon(k)b_{k}^{\dagger}b_{k}\\ \end{split}$ (S3)
Figure S1: The energy spectrum of free bosons with different $t_{2}/t_{1}$.
@@
Figure S2: The energy difference of the FDBH model with respect to the next-nearest-neighbor hopping. The red squares represent the energy difference between the ground state and the first excited state with $S_{\text{total}}^{z}=0$. The blue circles represent the energy difference between the ground state of $S_{\text{total}}^{z}=1$ and $S_{\text{total}}^{z}=0$. For $L=12$, the crossing point occurs at $t_{2c}$=0.37. For $L=20$, the crossing point locates at $t_{2c}=0.39$.
B: Level crossing for the BKT transition
@@
To identify the phase transition via level crossings, we investigated the energy gap $\Delta E=E_{m}^{(n)}-E_{m^{\prime}}^{(n^{\prime})}$ of the FDBH model using exact diagonalization with lattice sizes $L=12$ and $L=20$ at $V=0.1$ for $0.50<t_{2}/t_{1}<$ 0.5 as illustrated in Fig.S2. Here, $E_{m}^{(n)}$ denotes energy levels where $m=0,\pm 1$ represents the total spin along the $z$-direction, and $n$ denotes the energy level (e.g., ground state, first excited state). The red squares depict the energy difference between the singlet ground state and the first excited state with $S_{\text{total}}^{z}=0$, which is non-degenerate except at the Majumdar-Ghosh point $t_{2}=0.5$. The blue circles correspond to the energy difference between ground states with $S_{\text{total}}^{z}=1$ and $S_{\text{total}}^{z}=0$. The critical point is identified where these two lines cross. For $L=12$, the XY phase to VBS phase transition occurs at $t_{2c}=0.37$, while for $L=20$, it shifts slightly to $t_{2c}=0.39$. This shift to higher $t_{2c}$ values with increasing lattice size indicates a pronounced finite-size effect. In the hard-core boson limit, spin conservation is related to particle number conservation through the equation,
$S_{\text{total}}^{z}=L/2-N_{\text{total}},$ (S9)
@@
Figure S3: The critical exponents of the correlation length for VBS-AFM transitions.
C: Correlation length critical exponent
@@
$\chi_{F}(\lambda_{m})\sim L^{2/\nu},$ (S10)
for second-order phase transitions. Here, $\nu$ denotes the correlation length critical exponent, and $\lambda_{m}$ represents the peak position of the fidelity susceptibility for a system of size $L$. The correlation length critical exponent is determined by fitting the logarithm of both sides of Eq. (S10):
$\ln\chi_{F}(\lambda_{m})\propto 2/\nu\ln L.$ (S11)
@@
In the vicinity of the phase transition, we keep the interaction strength fixed and compute the ground state fidelity susceptibility of the system across various sizes (including $L=128,144,192,240$, and $288$) as a function of the next-nearest-neighbor hopping $t_{2}/t_{1}$. By fitting the maximum fidelity against system size using the logarithmic relationship in Eq. (S11), we extract the slope of the fit, thereby determining the correlation length critical exponent, as illustrated in Fig. S3. From the figure, it is observed that the critical exponents decrease gradually with increasing interaction strength. This suggests that the system exhibits varying universality classes for different quantum critical points.

### v3

@@ IV Deconfined quantum critical point
where $\chi_{L}^{m}$ represents the fidelity susceptibility per site at the peak position $\lambda=\lambda_{m}$, which converges towards the critical point $\lambda_{c}$ as $L$ tends to infinity. As the AFM phase breaks the $\mathbb{Z}_{2}$ spin-flip symmetry and the VBS phase breaks translation symmetry, the direct continuous second-order phase transition occurring between the AFM phase and the VBS state indicates that the phase transition is a DQCP beyond the LGW theory. The continuous phase transition is further confirmed by the divergence of the correlation length $\xi$ with the increasing number of states $M$ [cf. Fig.5(e)], as computed using TenPy [69]. Additionally, the central charge $c=1$ is derived from the scaling of the entanglement entropy $S$ [cf. Fig.5(f)], using the relation $S=\frac{c}{6}\log(\xi)$ [13, 33, 34, 35, 38]. We note that the results obtained from iTEBD method are consistent with those from finite-size scaling (see also Appendices D and E for details).
Interestingly, the DQCP arises from the merging of two BKT transitions [33, 34, 35]. These transitions can be understood through the concept of domain walls, which are topological defects in one-dimensional systems. Beginning from the VBS phase (i.e at $t_{2}/t_{1}=0.7$ and $V=0$), increasing the interaction strength $V/t_{1}$ induces the formation of AFM domain walls. Domain walls in VBS bind the AFM phase (and vice versa), leading to the destruction of the VBS phase and the emergence of the AFM phase as $V/t_{1}$ increases further. Consequently, a direct transition between the VBS phase to the AFM phase occurs under conditions of large $t_{2}/t_{1}$ and $V/t_{1}$, indicating the presence of a DQCP. This DQCP is linked to the two BKT transitions through a multicritical point characterized by an emergent high symmetry.
@@
V Conclusion
In summary, we employed the DMRG and iTEBD methods to investigate novel physical phenomena in the frustrated dipolar Bose-Hubbard model and to provide a comprehensive phase diagram of the model through the exploration of fidelity susceptibilities and order parameters. We identified two gapped phases, the AFM phase and the VBS phase, along with one gapless XY phase, each exhibiting two BKT transitions and a DQCP. Remarkably, we found that the DQCP arises from the fusion of two BKT transitions in this model. Moreover, we propose that this model could be realized using dipolar bosons.
Acknowledgements.
G.S. thanks Luis Santos for the valuable comments. G.S. is appreciative of support from the NSFC under the Grants No. 11704186, "the Fundamental Research Funds for the Central Universities, NO. NS2023055". W.-L.Y is supported by the NSFC under Grant No. 12174194, Opening Fund of the Key Laboratory of Aerospace Information Materials and Physics (Nanjing University of Aeronautics and Astronautics), MIIT, Top-notch Academic Programs Project of Jiangsu Higher Education Institutions (TAPP), and stable supports for basic institute research under Grant No. 190101. Y.-N.W and W.-Y.Z are supported by the Postgraduate Research & Practice Innovation Program of Jiangsu Province under Nos. KYCX24_0526 and KYCX23_0347, respectively. S.-P. K is appreciative of support from the NSFC under the Grant Nos. 11974053 and 12174030, and the National Key R&D Program of China under Grant No. 2023YFA1406704. This work is partially supported by the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics.
Appendix A Non-interacting energy spectrum
@@
$\begin{split}b_{r}=\frac{1}{\sqrt{N}}\sum e^{ikj}b_{k},\\ b_{r}^{\dagger}=\frac{1}{\sqrt{N}}\sum e^{-ik^{\prime}j}b_{k^{\prime}}^{\dagger},\end{split}$ (11)
the model described in Eq.(10) can be transformed into momentum space,
$\begin{split}H&=t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{2ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-2ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}\\ &=\sum_{k}\left[t_{1}(e^{ik}+e^{-ik})+t_{2}(e^{2ik}+e^{-2ik})\right]b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\left(2t_{1}\cos{k}+2t_{2}\cos{2k}\right)b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\epsilon(k)b_{k}^{\dagger}b_{k}\\ \end{split}$ (12)
Figure A1: The energy spectrum of free bosons with different $t_{2}/t_{1}$.
@@
Appendix B Level crossing for the BKT transition
Figure B1: The energy difference of the FDBH model with respect to the next-nearest-neighbor hopping. The red squares represent the energy difference between the ground state and the first excited state with $S_{\text{total}}^{z}=0$. The blue circles represent the energy difference between the ground state of $S_{\text{total}}^{z}=1$ and $S_{\text{total}}^{z}=0$. For $L=12$, the crossing point occurs at $t_{2c}$=0.37. For $L=20$, the crossing point locates at $t_{2c}=0.39$.
@@ Appendix B Level crossing for the BKT transition
To identify the phase transition via level crossings, we investigated the energy gap $\Delta E=E_{m}^{(n)}-E_{m^{\prime}}^{(n^{\prime})}$ of the FDBH model using exact diagonalization with lattice sizes $L=12$ and $L=20$ at $V=0.1$ for $0.50<t_{2}/t_{1}<$ 0.5 as illustrated in Fig.B1. Here, $E_{m}^{(n)}$ denotes energy levels where $m=0,\pm 1$ represents the total spin along the $z$-direction, and $n$ denotes the energy level (e.g., ground state, first excited state). The red squares depict the energy difference between the singlet ground state and the first excited state with $S_{\text{total}}^{z}=0$, which is non-degenerate except at the Majumdar-Ghosh point $t_{2}=0.5$. The blue circles correspond to the energy difference between ground states with $S_{\text{total}}^{z}=1$ and $S_{\text{total}}^{z}=0$. The critical point is identified where these two lines cross. For $L=12$, the XY phase to VBS phase transition occurs at $t_{2c}=0.37$, while for $L=20$, it shifts slightly to $t_{2c}=0.39$. This shift to higher $t_{2c}$ values with increasing lattice size indicates a pronounced finite-size effect. In the hard-core boson limit, spin conservation is related to particle number conservation through the equation,
$S_{\text{total}}^{z}=L/2-N_{\text{total}},$ (18)
@@
Appendix C Correlation length critical exponent
Figure C1: The critical exponents of the correlation length for VBS-AFM transitions.
@@ Appendix C Correlation length critical exponent
$\chi_{F}(\lambda_{m})\sim L^{2/\nu},$ (19)
for second-order phase transitions. Here, $\nu$ denotes the correlation length critical exponent, and $\lambda_{m}$ represents the peak position of the fidelity susceptibility for a system of size $L$. The correlation length critical exponent is determined by fitting the logarithm of both sides of Eq. (19):
$\ln\chi_{F}(\lambda_{m})\propto 2/\nu\ln L.$ (20)
@@
In the vicinity of the phase transition, we keep the interaction strength fixed and compute the ground state fidelity susceptibility of the system across various sizes (including $L=128,144,192,240$, and $288$) as a function of the next-nearest-neighbor hopping $t_{2}/t_{1}$. By fitting the maximum fidelity against system size using the logarithmic relationship in Eq. (20), we extract the slope of the fit, thereby determining the correlation length critical exponent, as illustrated in Fig. C1. From the figure, it is observed that the critical exponents decrease gradually with increasing interaction strength. This suggests that the system exhibits varying universality classes for different quantum critical points.
Appendix D Correlation functions
Figure D1: Order parameters and correlation functions. (a) XY to VBS phase transition at $V/t_{1}=0.1$ as the function of $t_{2}/t_{1}$, (b) XY to AFM phase transition at $t_{2}/t_{1}=0.1$ as the function of $V/t_{1}$, (c) VBS to AFM phase transition at $t_{2}/t_{1}=0.7$ as the function of $V/t_{1}$, (d) Correlation functions for XY phase at $t_{2}/t_{1}=0.1,V/t_{1}=0.5$ for $L=96$ in PBCs, (e) Correlation functions for AFM phase at $t_{2}/t_{1}=0.7,V/t_{1}=2.7$ for $L=96$ in PBCs, (f) Correlation functions for VBS phase with $t_{2}/t_{1}=0.7,V/t_{1}=0.1$ for $L=192$ in OBCs.
In this section, we calculated the long-range correlation functions defined as,
$C^{\alpha}_{r}=\langle\sigma^{\alpha}_{1}\sigma^{\alpha}_{r+1}\rangle,$ (21)
with $\alpha=x,y,z$ for each phase in order to demonstrate the characteristics of each phase. Additionally, we compute the order parameters for finite size systems. The order parameters $O_{\text{VBS}}$ for the VBS phase is computed from Eq.(7), while the order parameters $O_{\text{AFM}}$ for the AFM phase can be derived in terms of correlation functions as [72, 73],
$C_{\text{AFM}}=\lim_{r\rightarrow\infty}C^{z}_{r}\rightarrow O_{\text{AFM}}^{2}.$ (22)
In the following, I will present the order parameters $C_{\text{AFM}}$ based on above correlation functions. In our calculations, $C_{\text{AFM}}$ is computed using $r=L/2$ for system sizes up to $L=128$ in periodic boundary conditions (PBCs), while the order parameters $O_{\text{VBS}}$ is obtained for system sizes up to $L=288$ under OBCs. The results for $C_{\text{AFM}}$ and $O_{\text{VBS}}$ are presented in Fig.D1(a-c), where it can be observed that they are consistent with the iTEBD results [c.f. Fig.3, Fig.4 and Fig.5]. The nature of the XY, AFM and VBS phases is further confirmed by the long-range correlation functions, as demonstrated in Fig.D1(d-f).
Appendix E DMRG simulations
In this section, we discuss the convergence of the ground state energy of the system. During the DMRG simulations, we gradually increase the number of states to a maximum of $M=500$ for different lattice sizes, as well as the number of DMRG sweeps to a maximum of 40 until the system converged. We benchmark our results by comparing the energy differences of the system at $t_{2}/t_{1}=0.7$ for various states as shown in Fig.E1. We find that under OBCs, the energy difference between $M=400$ and $M=500$ states for $L=288$ is very small, reaching $10^{-9}$, while under PBCs, the energy difference between $M=400$ and $M=500$ states for $L=96$ is also very small, reaching $10^{-5}$. The results from the finite-size DMRG simulations are in good agreement with the iTEBD results.
Figure E1: The difference in ground state energy. (a) the energy difference $\Delta E_{0}$ for $L=288$ at $t_{2}/t_{1}=0.7$ under OBCs, (b) the energy difference $\Delta E_{0}$ for $L=96$ at $t_{2}/t_{1}=0.7$ under PBCs.